### v3

# Stanford CA 94305, USA

Abstract
Quasi-Monte Carlo sampling can attain far better accuracy than plain Monte Carlo sampling. However, with plain Monte Carlo sampling it is much easier to estimate the attained accuracy. This article describes methods old and new to quantify the error in quasi-Monte Carlo estimates. An important challenge in this setting is that the goal of getting accuracy conflicts with that of estimating the attained accuracy. A related challenge is that rigorous uncertainty quantifications can be extremely conservative. A recent surprise is that some RQMC estimates have nearly symmetric distributions and that has the potential to allow confidence intervals that do not require either a central limit theorem or a consistent variance estimate.
keywords:
Bootstrap, Digital shifts, Median of means, Randomized quasi-Monte Carlo, Scrambling
1 Introduction
Quasi-Monte Carlo (QMC) sampling is used to compute numerical approximations to integrals. It is an alternative to plain Monte Carlo (MC) sampling. Under reasonable conditions, QMC attains far better accuracy, at least asymptotically, than MC does. MC retains one advantage: it is easier to estimate the attained accuracy of an MC estimate than it is for a QMC estimate. We will see below how randomized QMC (RQMC) supports error estimation, but even there, plain MC makes the task easier.
Let $\mu$ be the integral of interest and $\hat{\mu}_{n}$ be our estimate of $\mu$ using $n$ evaluations of the integrand $f$. We want $\varepsilon_{n}=|\hat{\mu}_{n}-\mu|$ to be small but we also want to know something about how large this error is, and of course we cannot use $\mu$ to do that. Estimating $\hat{\mu}-\mu$ or a distribution for it is a form of uncertainty quantification (UQ). In a setting with high stakes (think of safety-related problems), we will want to have high confidence that $\mu$ is in some interval $[a,b]$ all points of which are acceptable. Then it is not enough for $|\hat{\mu}_{n}-\mu|$ to be small. It should also be known to be small. For some problems we might only need to verify that $a\leqslant\mu$ or that $\mu\leqslant b$.
An ideal form of UQ arises in the form of a ‘certificate’: two computable numbers $a\leqslant b$ where we are mathematically certain that $a\leqslant\mu\leqslant b$. Equivalently, having an estimate $\hat{\mu}_{n}$ with certainty that $|\hat{\mu}_{n}-\mu|\leqslant\Delta$ for known $\Delta<\infty$ provides a certificate. These are rare for integration problems, but we will see a few examples. It is much more common for our theorems to provide bounds for $|\hat{\mu}-\mu|$ that we cannot compute. Existence of a finite error bound does not meaningfully quantify uncertainty.
We will look at several approaches to UQ that all require some extra knowledge. One approach is to suppose that $f$ belongs to a given function class ${\mathcal{F}}$, in which we can bound $\varepsilon_{n}=\varepsilon_{n}(f)$. Another is to use a model with random $\hat{\mu}_{n}$ where we can get bounds on the distribution of $\varepsilon_{n}$. Sometimes we can only get asymptotically justified bounds as the amount of computation grows to infinity.
We will assume that the method for getting a confidence interval involves randomizing the points at which $f$ is evaluated, via a pseudo-random number generator. This depends on using a well tested pseudo-random number generator. The big crush of [63] provides such a thorough test. It is also possible to use a Bayesian model in which $f$ itself is random. The most common choice is a Gaussian process model for $f$. For an entry point to that large literature, one can start with [9]. The Bayesian approach can produce very accurate estimates but there is no counterpart to the big crush for that approach, and so it does not as yet have a well tested objective way to estimate error.
Methods that estimate both $\mu$ and $\varepsilon_{n}$ have a tradeoff to make. This is most clear in RQMC (see below) which uses $R$ random replicates of an integration rule on $n$ function evaluations. For a given cost $nR$ we get more accuracy in $\hat{\mu}$ by taking larger $n$ and better UQ by taking larger $R$ (hence smaller $n$).
A second tradeoff can arise. We may have certainty or just high confidence that $-\Delta_{n}\leqslant\hat{\mu}_{n}-\mu\leqslant\Delta_{n}$ where we also know that $|\hat{\mu}_{n}-\mu|/\Delta_{n}$ converges to zero, either deterministically or probabilistically. Then our resulting estimate $\Delta_{n}$ of the size of the error will for large enough $n$ be a significant exaggeration, by an unknown factor. The bounds can be very conservative.
Since this article is to appear in an MCQMC proceedings, we assume some background knowledge that would take too much space to include here. For readers new to those topics, here are some bibliographic references that also include some history. Those topics include discrepancy [13], the ANOVA decomposition [79, Appendix A], weighted Hilbert spaces [27], digital QMC constructions [28] (such as Sobol’ sequences and Faure sequences) integration lattices [92] and their randomizations [57]. Some further references are given in context. Further background is given on confidence interval methods as those are not widely studied in the QMC literature.
The contents of this paper are as follows. Section 2 introduces some notation along with basic strategies to get an estimate $\hat{\mu}$ of $\mu$. Section 3 describes the basics of uncertainty quantification. There are certificates, confidence intervals, and asymptotic confidence intervals. The results making the more desirable claims are harder to apply. Section 4 gives some classical certificates (bracketing inequalities) for numerical integration of convex functions along with a new one for completely monotone integrands using QMC points with nonnegative and nonpositive local discrepancy. Section 5 covers RQMC. It emphasizes asymptotic statistical confidence intervals as $n\to\infty$ and how hard they are to apply to RQMC with a small number $R$ of independent replicates of an RQMC rule on $n$ points. An empirical investigation in [61] showed that classical $t$-test intervals based on $R$ independent replicates are surprisingly reliable and there are now a few theoretical explanations to show why. Section 6 looks at a proposal by Tony Warnock and John Halton that in the notation of this paper is like using $R$ quasi-random replicates of a QMC rule instead of $R$ random replicates. Their quasi-standard error has had some empirical success and also has some failings. The idea might benefit from a further investigation. Section 7 describes the guaranteed automatic integration library which aims to sample adaptively until a specified error criterion $|\hat{\mu}-\mu|\leqslant\epsilon$ is met, either absolutely or probabilistically. Section 8 presents areas of current active interest in UQ for QMC: unbiased MCMC, normalizing flows, median of means and results about $R$ growing with $n$. Conclusions in Section 9 describe how the present understanding of the problem has many gaps. It appears that there will be a role for the analysis of distributions that become symmetric as $n\to\infty$ without becoming Gaussian.
2 Notation and basic estimates of $\mu$
For integers $d\geqslant 1$, we use $1{:}d$ to represent the set $\{1,2,\dots,d\}$. The cardinality of $u\subseteq 1{:}d$ is denoted by $|u|$ and we use $-u=1{:}d\setminus u$. For ${\bm{x}}\in[0,1]^{d}$ and $u\subseteq 1{:}d$ we use ${\bm{x}}_{u}\in[0,1]^{|u|}$ to represent the components of ${\bm{x}}$ with indices in $u$. For ${\bm{x}},{\bm{z}}\in[0,1]^{d}$ we write ${\bm{x}}_{u}{:}{\bm{z}}_{-u}$ for the point with $j$’th component $x_{j}$ when $j\in u$ and $j$’th component $z_{j}$ when $j\not\in u$. We abbreviate ${\bm{x}}_{-\{j\}}{:}{\bm{z}}_{\{j\}}$ to ${\bm{x}}_{-j}{:}{\bm{z}}_{j}$. We use ${\bm{1}}=(1,1,\dots,1)\in{\mathbb{R}}^{d}$ and ${\bm{0}}=(0,0,\dots,0)\in{\mathbb{R}}^{d}$. For $u\subseteq 1{:}d$, we use $\partial^{u}f$ for the partial derivative of $f$ taken once with respect to each $x_{j}$ for $j\in u$. By convention $\partial^{\varnothing}f=f$. Our estimate for the integral will be denoted by $\hat{\mu}$ or by $\hat{\mu}_{n}$ when it is important to emphasize the sample size $n$. When we have independent replicates, the $r$’th one will be denoted by $\hat{\mu}^{(r)}$ or $\hat{\mu}_{n}^{(r)}$ as needed.
We say that $a_{n}=o(b_{n})$ if $a_{n}/b_{n}\to 0$ as $n\to\infty$. We say that $a_{n}=O(b_{n})$ if there is finite $c$ with $|a_{n}|\leqslant cb_{n}$ for all but finitely many integers $n\geqslant 1$. We say that $a_{n}=\Omega(b_{n})$ if there is $c>0$ with $a_{n}\geqslant cb_{n}$ for all but finitely many $n$. We say that $a_{n}=\Theta(b_{n})$ if $a_{n}=O(b_{n})$ and $a_{n}=\Omega(b_{n})$.
This paper surveys many different results from multiple literatures. As a result, it is necessary for some symbols to be used with more than one meaning, even as some symbols common in the literature have been changed to reduce the number of conflicts here. The specific use cases are different enough to be clear from context.
We focus on the problem of approximating
$\mu=\int_{[0,1]^{d}}f({\bm{x}})\,{\mathrm{d}}{\bm{x}}$
for $d\geqslant 1$. Formulating a given problem this way may require extensive use of transformations such as those in [21] in order to handle domains other than $[0,1]^{d}$ and distributions other than the uniform one. Section 8 points to some new work using normalizing flows to augment those transformations. For very smooth $f$ and small $d$, classical integration methods from [19] are quite effective at estimating $\mu$.
For larger $d$ and/or less regular $f$, MC methods can be more accurate. In plain MC we take ${\bm{x}}_{i}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}{\mathbb{U}}[0,1]^{d}$ and then
$\hat{\mu}_{n}=\frac{1}{n}\sum_{i=1}^{n}f({\bm{x}}_{i}).$ (1)
If $\mu$ exists then $\Pr(\lim_{n\to\infty}\hat{\mu}_{n}=\mu)=1$ by the strong law of large numbers. If
$\sigma^{2}=\int_{[0,1]^{d}}(f({\bm{x}})-\mu)^{2}\,{\mathrm{d}}{\bm{x}}<\infty$
then ${\mathbb{E}}((\hat{\mu}_{n}-\mu)^{2})=\sigma^{2}/n$ and the root mean squared error (RMSE) is $\sigma/\sqrt{n}$. The $n^{-1/2}$ rate is not as good as the ones for smooth low dimensional functions in [21] but remarkably, it requires no differentiability and is the same in any dimension.
In QMC, we replace uniform random ${\bm{x}}_{i}$ by carefully chosen deterministic points ${\bm{x}}_{i}\in[0,1]^{d}$ before estimating $\mu$ by $\hat{\mu}_{n}$ with the same simple average (1) that we use for MC. These points are usually chosen by methods that for large $n$, have a small value for the star discrepancy
$D_{n}^{*}$ $=D_{n}^{*}({\bm{x}}_{1},\dots,{\bm{x}}_{n})=\sup_{{\bm{a}}\in[0,1]^{d}}|\delta({\bm{a}})|,\quad\text{where}$
$\delta({\bm{a}})$ $=\frac{1}{n}\sum_{i=1}^{n}{\mathbbold{1}}\{{\bm{x}}_{i}\in[{\bm{0}},{\bm{a}})\}-\prod_{j=1}^{d}a_{j}$
is the local discrepancy at ${\bm{a}}$. The QMC counterpart to the law of large numbers is that if $f$ is Riemann integrable and $D_{n}^{*}\to 0$ as $n\to\infty$, then $\hat{\mu}_{n}\to\mu$. See [73] which also mentions a converse that if $\hat{\mu}_{n}\to\mu$ whenever $D_{n}^{*}\to 0$, then $f$ must be Riemann integrable. We will see below that if $f$ is of bounded variation in the sense of Hardy and Krause (written $f\in\mathrm{BVHK}$) then $\varepsilon_{n}=O(D_{n}^{*})$. Since it is possible to find points with $D_{n}^{*}=O(n^{-1+\epsilon})$ for any $\epsilon>0$, we $\hat{\mu}_{n}=O(n^{-1+\epsilon})$, and then we can get asymptotically lower errors from QMC than the RMSE of MC.
The $\epsilon$ in that rate for $\varepsilon_{n}$ hides powers of $\log(n)$. Those are present in some worst cases, but they do not seem to appear in applications [81]. If $f$ has a continuous mixed partial derivative taken once with respect to each of the $d$ components of ${\bm{x}}$, then we can show that $f\in\mathrm{BVHK}$ using a result from [37]. That dominating mixed partial derivative falls short of the complete $r$-fold differentiability that appears in the $O(n^{-r/d})$ result mentioned earlier, and the price we pay is in logarithmic factors.
We will also consider RQMC, a hybrid of MC and QMC. In RQMC, the points ${\bm{x}}_{1},\dots,{\bm{x}}_{n}$ are chosen so that individually ${\bm{x}}_{i}\sim{\mathbb{U}}[0,1]^{d}$ while collectively these points have low discrepancy: for any $\epsilon>0$ and $B>0$, $\Pr(D_{n}^{*}({\bm{x}}_{1},\dots,{\bm{x}}_{n})\leqslant Bn^{-1+\epsilon})=1$ holds for all but finitely many integers $n\geqslant 1$. We write this as $D_{n}^{*}=O(n^{-1+\epsilon})$. For a survey of RQMC, see [58].
3 Uncertainty quantification
Uncertainty quantification takes things to the next level, where we want to know something about $\mu_{n}-\mu$ or $|\mu_{n}-\mu|$. In this section we review some well known UQ methods for integration methods. Perfect knowledge of $\mu_{n}-\mu$ implies perfect knowledge of $\mu$ and then there is no more uncertainty to quantify. A known distribution for $\mu_{n}-\mu$ does not collapse that way. Neither does perfect knowledge of $|\hat{\mu}_{n}-\mu|$; it simply means that $\mu=\hat{\mu}_{n}\pm|\hat{\mu}_{n}-\mu|$. Some methods provide information about the distribution of $|\hat{\mu}_{n}-\mu|$. We start with the more desirable UQ conclusions and then introduce some compromises in order to get more usable methods.
By the Koksma-Hlawka inequality [49],
$\varepsilon_{n}\leqslant D_{n}^{*}({\bm{x}}_{1},\dots,{\bm{x}}_{n})V_{\mathrm{HK}}(f)$ (2)
where $V_{\mathrm{HK}}(f)$ is the total variation of $f$ in the sense of Hardy and Krause. See [77] for a definition and some properties of $V_{\mathrm{HK}}$. This bound has absolute certainty, it applies to the specific integrand we have, the value of $n$ that we used, and even the precise list of points ${\bm{x}}_{i}$ that we used.
Equation (2) falls short of providing an uncertainty quantification because we ordinarily do not know either of the factors in it. It is expensive to compute $D_{n}^{*}$. The commonly used algorithm of [30] costs $O(n^{1+d/2})$. Let us suppose that it costs $\Omega(n^{1+d/2})$. Then in the time it would take us to compute $D_{n}^{*}$ for QMC we could get $N=\Omega(n^{1+d/2})$ points in MC. Then MC would get an RMSE of $O(n^{-1/2-d/4})$ in the time that QMC using known $D_{n}^{*}$ would get an error of $O(n^{-1+\epsilon})$. We could in principle precompute $D_{n}^{*}$ for a number of point sets, but then we would face the second problem: $V_{\mathrm{HK}}(f)$ is extremely hard to compute, far worse than $\mu$ itself. For smooth enough $f$, it is a sum of $12^{d}-$ 1 integrals of the absolute values of some mixed partial derivatives of $f$ [77]. For less smooth $f$, the problem is even harder. As a result, (2) serves to show us that QMC can be much better than MC, but it does not generally let us verify that that has happened for given $f$ and $n$.
More modern bounds are derived for reproducing kernel Hilbert spaces (RKHSs) of integrands. The unanchored Sobolev spaces described in [27] are a prominent example. For every $u\subseteq 1{:}d$, define a weight $\gamma_{u}>0$ and then let ${\bm{\gamma}}$ represent all $2^{d}$ of those weights. The squared norm in this space is
$\|f\|_{{\bm{\gamma}}}^{2}=\sum_{u\subseteq 1{:}d}\frac{1}{\gamma_{u}}\int_{[0,1]^{|u|}}\biggl{(}\int_{[0,1]^{d-|u|}}\partial^{u}f({\bm{x}}){\mathrm{d}}{\bm{x}}_{-u}\biggr{)}^{2}{\mathrm{d}}{\bm{x}}_{u},$ (3)
for weak partial derivatives $\partial^{u}f$. The error $|\hat{\mu}-\mu|$ is upper bounded by $\|f\|_{{\bm{\gamma}}}$ times a corresponding $L_{2}$ discrepancy measure defined by a reproducing kernel [27]. The widely used product weights have $\gamma_{u}=\prod_{j\in u}\gamma_{j}$ for $\gamma_{j}>0$. When $\gamma_{j}$ decays rapidly with increasing $j$, for example with $\gamma_{j}=j^{-\eta}$ for $\eta>1$, then strong tractability (see [93]) holds, in which the number of function evaluations required to get an error below $\epsilon\|f\|_{{\bm{\gamma}}}$ does not depend on the dimension $d$. For $\eta>2$, strong tractability holds with errors $O(n^{-1+\delta})$ for any $\delta>0$. This tractability property has been widely used to design better QMC point sets. However evaluating $\|f\|_{{\bm{\gamma}}}$ for an integrand $f$ is usually quite difficult and so, just like the Koksma-Hlawka inequality, the error bounds in weighted spaces do not provide an uncertainty quantification.
The MC setting is more favorable for UQ. By the Chebyshev inequality we know that for $\lambda>1$,
$\Pr\Bigl{(}|\hat{\mu}_{n}-\mu|\geqslant\frac{\lambda\sigma}{\sqrt{n}}\Bigr{)}\leqslant\frac{1}{\lambda^{2}}.$ (4)
Equation (4) is only a probabilistic bound on $\varepsilon_{n}$, rather than a certain bound like (2). However, the unknown quantity $\sigma$ in it is much easier to work with than $V_{\mathrm{HK}}$ or $\|f\|_{{\bm{\gamma}}}$ is. In some settings we may know $\sigma$, or a bound for $\sigma$. For example, if $0\leqslant f({\bm{x}})\leqslant 1$, then $\sigma\leqslant 1/2$.
From the Chebyshev inequality we see that
$\Pr\Bigl{(}\hat{\mu}_{n}-\frac{\lambda\sigma}{\sqrt{n}}\leqslant\mu\leqslant\hat{\mu}_{n}+\frac{\lambda\sigma}{\sqrt{n}}\Bigr{)}\geqslant 1-\frac{1}{\lambda^{2}}.$ (5)
For known $\sigma$, equation (5) describes a confidence interval: a computable random interval with at least the desired probability of containing the true value of $\mu$. To have that probability be at least $99$% we may take $\lambda=10$. We have given up the certainty of our bound in order to get a method based on $\sigma$ instead of $V_{\mathrm{HK}}$ or $\|f\|_{{\bm{\gamma}}}$. The Chebyshev inequality (4) is tight because for any $\lambda>1$ there is a distribution for $\varepsilon_{n}$ that makes it an equality.
The Chebyshev confidence interval is not commonly used. Even when we have a bound for $\sigma$, the interval can be very conservative. As mentioned above, to get 99% confidence, we can use $\lambda=10$. A Gaussian random variable only has probability about $1.5\times 10^{-23}$ to be $10$ or more standard deviations from its mean. Here the error probability bound of $0.01$ is quite loose, despite the fact that it came from an inequality that is tight.
The usual UQ for MC is based on the central limit theorem (CLT) under which
$\lim_{n\to\infty}\Pr\Bigl{(}\frac{\hat{\mu}_{n}-\mu}{\sigma/\sqrt{n}}\leqslant z\Bigr{)}=\Phi(z)$ (6)
where $\Phi$ is the cumulative distribution function of the Gaussian distribution with mean 00 and variance $1$. In addition to being probabilistic, it is also asymptotic in $n$. Its popularity stems from its wide applicability after we estimate $\sigma$ from the same data we use for $\hat{\mu}$.
The CLT typically gives much narrower confidence intervals than Chebyshev’s inequality provides. In plain MC with $n\geqslant 2$ we can estimate $\sigma^{2}$ by
$s^{2}=\frac{1}{n-1}\sum_{i=1}^{n}(f({\bm{x}}_{i})-\mu)^{2}.$ (7)
The peculiar denominator $n-1$ makes ${\mathbb{E}}(s^{2})=\sigma^{2}$. If $\sigma^{2}<\infty$, then equation (6) also holds with $s$ replacing $\sigma$. Then for $10<\alpha<$ 1,
$\lim_{n\to\infty}\Pr\Bigl{(}\hat{\mu}_{n}-\frac{s}{\sqrt{n}}\Phi^{-1}(1-\alpha/2)\leqslant\mu\leqslant\hat{\mu}_{n}+\frac{s}{\sqrt{n}}\Phi^{-1}(1-\alpha/2)\Bigr{)}=1-\alpha.$ (8)
Commonly chosen values for $\alpha$ are $0.05$ and $0.01$ which provide, respectively, asymptotic $95$% and $99$% confidence intervals for $\mu$. Here $\Phi^{-1}(0.975)\doteq 1.96$ and $\Phi^{-1}(0.995)\doteq 2.58$.
The usual statistical confidence interval is a bit wider than the one in (8) because it makes an adjustment for the sampling uncertainty in $s^{2}$. The interval in equation (8) is not exact, even when $f({\bm{x}})$ has a Gaussian distribution. We may correct this using
$\lim_{n\to\infty}\Pr\Bigl{(}\hat{\mu}_{n}-\frac{s}{\sqrt{n}}t^{1-\alpha/2}_{(n-1)}\leqslant\mu\leqslant\hat{\mu}_{n}+\frac{s}{\sqrt{n}}t^{1-\alpha/2}_{(n-1)}\Bigr{)}=1-\alpha$ (9)
where $t^{1-\alpha/2}_{(n-1)}$ is the $21-\alpha/$ 2 quantile of Student’s $t$ distribution on $n-1$ degrees of freedom. The correction for degrees of freedom makes a meaningful difference when $n$ is small which is valuable for RQMC as we will see in Section 5 . The coverage is exactly $1-\alpha$ if $f({\bm{x}}_{i})$ are Gaussian (and $n\geqslant 2$) and the limit (9) holds if $f({\bm{x}})$ has finite variance.
We would prefer a non-asymptotic and nonparametric confidence interval. That would provide exactly $1-\alpha$ coverage probability without requiring that $f({\bm{x}})$ have a distribution belonging to a finite dimensional family, such as the Gaussian distributions. Unfortunately, exact nonparametric confidence intervals do not exist under conditions given in Bahadur and Savage [5]. They consider a set ${\mathcal{F}}$ of distributions on ${\mathbb{R}}$. Letting $\mu(F)$ be ${\mathbb{E}}(Y)$ when $Y\sim F\in{\mathbb{R}}$, their conditions are:
  1.  (i)
    For all $F\in{\mathcal{F}}$, $\mu(F)$ exists and is finite.
  2.  (ii)
    For all $m\in{\mathbb{R}}$ there is $F\in{\mathcal{F}}$ with $\mu(F)=m$.
  3.  (iii)
    ${\mathcal{F}}$ is convex: if $F,G\in{\mathcal{F}}$ and $10<\pi<$ 1, then $\pi F+(1-\pi)G\in{\mathcal{F}}$.
Under these conditions, their Corollary 2 shows that a Borel set constructed based on $Y_{1},\dots,Y_{N}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}F$ that contains $\mu(F)$ with probability at least $1-\alpha$ also contains any other $m\in{\mathbb{R}}$ with probability at least $1-\alpha$. More precisely: we can get a confidence set, but not a useful one. They allow $N$ to be random so long as $\Pr(N<\infty)=1$.
There are settings where non-asymptotic confidence intervals can be constructed. To get them, we need some extra knowledge about the distribution of $f({\bm{x}})$ for ${\bm{x}}\sim{\mathbb{U}}[0,1]^{d}$ that does not come from sampling. We saw above that knowing $\sigma$ allows us to get a conservative confidence interval based on the Chebyshev inequality.
Suppose that independent $f({\bm{x}}_{i})$ satisfies known finite upper and lower bounds. Hoeffding’s inequality [52] then provides confidence intervals for $\mu$ in terms of those bounds. In the MC context it is natural to assume that the bounds are the same for all $i=1,\dots,n$. Because the bounds are known we can make a linear adjustment to $f$ and then without loss of generality we may then suppose that $0\leqslant f({\bm{x}})\leqslant 1$ for all ${\bm{x}}\in[0,1]^{d}$. Then for $t>0$, Hoeffding’s inequality gives
$\Pr(|\hat{\mu}-\mu|\geqslant t/n)\leqslant 2\exp(-t^{2}/n).$ (10)
This holds for any distribution of $f({\bm{x}})$ with $\Pr(0\leqslant f({\bm{x}})\leqslant 1)=1$. It is possible because that set of bounded random variables does not satisfy Bahadur and Savage’s condition (ii).
A very natural setting with known bounds arises when $f({\bm{x}})$ takes the value $1$ for ${\bm{x}}\in A$ and 00 otherwise. Then $\mu$ can be interpreted as the probability that ${\bm{x}}\in A$ when ${\bm{x}}\sim{\mathbb{U}}[0,1]^{d}$.
Because (10) holds for any random variable bounded between 00 and $1$, it is very conservative for some of them. For instance if $|f({\bm{x}})-1/2|\leqslant 1/2000$ always holds then the Hoeffding interval is 1000 times wider than it has to be. This extra width is a consequence of imperfect knowledge that specifies a wider than necessary interval. Empirical Bernstein inequalities [70] make use of the sample variance from (7) to get less conservative intervals that nonetheless always have the desired coverage level.
There is ongoing work in forming confidence intervals that have guaranteed coverage for bounded random variables and finite $n$ while being asymptotically as narrow as possible. See [4] and [101]. Some of this work allows for dependence among the values of $f({\bm{x}}_{i})$. Some of it provides confidence statements that are always valid in that they produce intervals $[a_{n},b_{n}]$ with $\Pr(a_{n}\leqslant\mu\leqslant b_{n},\forall n\geqslant 1)\geqslant 1-\alpha$. These confidence sequences have to be wider at each $n$ than if we only require validity at one value of $n$.
When $0\leqslant f({\bm{x}})\leqslant 1$ holds then this also holds for an average of $n$ evaluations of $f$ at some RQMC points. Jain et al. [53] study confidence interval widths using the empirical Bernstein and related betting methods from [101] for RQMC using $R$ random replicates of $n$ RQMC points. They impose a budget constraint that $nR=N$. If $f$ is regular enough for RQMC to be very effective, then the optimal $n$ grows only very slowly as $N\to\infty$. The resulting confidence intervals narrow at a faster rate in $N$ than the ones based on Monte Carlo do, but they narrow at a slower rate in $N$ than the RQMC standard deviations do.
4 Certificates and bracketing
A certificate is a (finite) computable number that is mathematically guaranteed to be no smaller than $\varepsilon_{n}$. This concept is widely used in convex optimization [8]. There the goal is to minimize a convex function $f$ over ${\bm{x}}\in C$ for a convex set $C$. Once $f$ has been evaluated at points ${\bm{x}}_{i}$ for $i=1,\dots,n$, we can be sure that the minimum is no larger than $\min(f({\bm{x}}_{1}),\dots,f({\bm{x}}_{n}))$. This upper bound does not require convexity of $f$. If $f$ is convex, then with enough data, we can get a lower bound. A convex function cannot go below any of its supporting hyperplanes. Then $f$ has to be above the pointwise maximum of all its supporting hyperplanes that the algorithm has encountered. Now the minimum value of that pointwise maximum provides a computable lower bound for the minimum of $f$. See Chapter 5 of [8] on duality. The user then has certainty that the answer lies between two computable numbers $a$ and $b$. If we take $(a+b)/2$ as the estimate then we know that the error is no larger than $(b-a)/2$.
There are a few instances in numerical integration where we can get a certificate. They are known as bracketing inequalities. Here we review those cases and then present a recent quasi-Monte Carlo version based on nonnegative local discrepancy and completely monotone functions.
It is obvious that if $f({\bm{x}})$ obeys an upper bound on a set, then so does its average over that set. Similarly for lower bounds. Let $f$ be a nondecreasing function of $x$ over $[0,1]$. Then $f((i-1)/n)\leqslant f(x)\leqslant f(i/n)$ for all $x\in[(i-1)/n,i/n]$ and $i=1,\dots,n$. This allows us to bracket $\mu$ between the values of the left and right endpoint rules:
$\hat{\mu}_{\mathrm{Left}}:=\frac{1}{n}\sum_{i=1}^{n}f\Bigl{(}\frac{i-1}{n}\Bigr{)}\leqslant\mu\leqslant\frac{1}{n}\sum_{i=1}^{n}f\Bigl{(}\frac{i}{n}\Bigr{)}=:\hat{\mu}_{\mathrm{Right}}.$ (11)
By raising the computation from $n$ to $n+1$ function evaluations we get an interval for $\mu$.
A more interesting bracketing inequality holds when $f$ is a convex function on $[0,1]$. Now the average of $f$ over $[(i-1)/n,i/n]$ is no smaller than $f((i-1/2)/n)$ by Jensen’s inequality and no larger than $f(((i-1)/n)+f(i/n))/2$ by a secant inequality. As a result, $\hat{\mu}_{\mathrm{Mid}}\leqslant\mu\leqslant\hat{\mu}_{\mathrm{Trap}}$ where
$\hat{\mu}_{\mathrm{Mid}}$ $=\frac{1}{n}\sum_{i=1}^{n}f\Bigl{(}\frac{i-1/2}{n}\Bigr{)}\quad\text{and}\quad\hat{\mu}_{\mathrm{Trap}}=\frac{1}{n}\Bigl{[}\frac{1}{2}f(0)+\sum_{i=1}^{n-1}f\Bigl{(}\frac{i}{n}\Bigr{)}+\frac{1}{2}f(1)\Bigr{]}$ (12)
are the midpoint and trapezoid rules, respectively. This bracketing requires $12n+$ 1 function evaluations because there is no overlap among the points used by the two rules.
For smooth enough integrands, these error bounds from bracketing inequalities are far wider than necessary [19, Chapter 2]. If $f^{\prime}$ is integrable, then the left and right endpoint rules have an error that is asymptotically $O(1/n)$. If we average them then we get the trapezoid rule with error $O(1/n^{2})$ if $f^{\prime}\in L_{2}[0,1]$ [18].
The same issue arises when we bracket the integrand between the midpoint and trapezoid rules. If $f^{\prime\prime}$ is continuous, then $\hat{\mu}_{\mathrm{Mid}}=\mu-f^{\prime\prime}(z_{\mathrm{M}})/(24n^{2})$ and $\hat{\mu}_{\mathrm{Trap}}=\mu+f^{\prime\prime}(z_{\mathrm{T}})/(12n^{2})$ for points $z_{\mathrm{M}},z_{\mathrm{T}}\in(0,1)$. These oppositely signed errors can be largely canceled by Simpson’s rule. It satisfies $\hat{\mu}_{\mathrm{Simp}}=(2\hat{\mu}_{\mathrm{Mid}}+\hat{\mu}_{\mathrm{Trap}})/3$. If $f^{(4)}$ is continuous on $[0,1]$ then by equation (2.2.6) of [19]
$\hat{\mu}_{\mathrm{Simp}}=\mu+\frac{f^{(4)}(z_{\mathrm{S}})}{180n^{4}}$
for some $z_{\mathrm{S}}\in(0,1)$. Note that this estimate uses $12n+$ 1 function evaluations. Then for a convex function with four continuous derivatives we get an estimate with error $O(n^{-4})$ known to lie within a computable interval of width $O(n^{-2})$. It is hard to just barely bound an error.
There is a discussion of practical error estimation in [19, Chapter 4.9]. Many of them require special conditions on $f$. One of the most applicable methods is to get two estimates, such as $\hat{\mu}_{n}$ and $\hat{\mu}_{2n}$. Then $|\hat{\mu}_{2n}-\hat{\mu}_{n}|$ is a rough estimate of the error in $\hat{\mu}_{n}$. It is also likely to be an overestimate of the error in $\hat{\mu}_{2n}$ which will usually be a better estimate than $\hat{\mu}_{n}$. Despite its wide use, this method can clearly give an unreliable estimate. There is an analysis in [68] showing how using the difference between two different integration rules can be an unreliable way to quantify uncertainty.
For $d=1$ and a function with $r\geqslant 1$ continuous derivatives we can get estimates of $\mu$ with an error of $O(n^{-r})$ with an implied constant that depends on $\|f^{(r)}\|_{\infty}$. For errors that decay like $n^{-r}$, an investigator may know enough about their integrand to reason that $\|f^{(r)}\|_{\infty}$ cannot be extremely large compared to $n^{r}$ times the implied constant in an error bound. Then, even lacking a known bound for $\|f^{(r)}\|_{\infty}$ they may be confident that some value of $n$ is good enough. For example, they might be confident that the error $|\hat{\mu}_{n}-\mu|$ is comparable in size to a floating point error that they already find acceptable.
The situation is very different for multivariate problems. Then there is a curse of dimension from [6]. Here is a sketch based on [29]. Suppose that the absolute value of the partial derivative of $f$ taken $a_{j}\geqslant 0$ times with respect to component $x_{j}$ is never larger than some $M\in(0,\infty)$ for any ${\bm{x}}\in[0,1]^{d}$ when $\sum_{j=1}^{d}a_{j}\leqslant r$. Then there still exists $B>0$ such that for any rule of the form $\sum_{i=1}^{n}w_{i}f({\bm{x}}_{i})$ for $w_{i}\in{\mathbb{R}}$ and ${\bm{x}}_{i}\in[0,1]^{d}$ there is $f$ with $|\hat{\mu}(f)-\mu(f)|>Bn^{-r/d}$. For randomized points ${\bm{x}}_{i}$, the RMSE cannot be below $B^{\prime}n^{-1/2-r/d}$ for some $B^{\prime}>0$. For large $d$, we cannot simply ignore the integration error. It is also not easy to use methods tuned for $r\geqslant 3$. Dimov [29] describes them as ‘sophisticated’.
Multivariate certificates for integration are rare. We can apply bounds (11) and (12) to iterated integrals. Then if $f$ is nondecreasing in each component of ${\bm{x}}$ or if $f$ is convex, we get certificates but only at rates $O(n^{-1/d})$ and $O(n^{-2/d})$ for estimates that have errors $O(n^{-2/d})$ and $O(n^{-4/d})$, respectively, when $f$ is smooth enough.
Another use of convexity is given in [43]. For a convex function $f$ defined on a simplex ${\mathbb{S}}\subset{\mathbb{R}}^{d}$, with vertices at ${\bm{x}}_{0},{\bm{x}}_{1},\dots,{\bm{x}}_{d}$
$f\Bigl{(}\frac{1}{d+1}\sum_{i=0}^{d}{\bm{x}}_{i}\Bigr{)}\leqslant\frac{1}{\mathrm{vol}(S)}\int_{S}f({\bm{x}}){\mathrm{d}}{\bm{x}}\leqslant\frac{1}{d+1}\sum_{i=0}^{d}f({\bm{x}}_{i}).$
This method can be applied to domains that are partitioned into simplices, allowing us to generalize the rule in (12) while reusing any function evaluations that are on a vertex of more than one simplex. For integration over $[0,1]^{d}$ partitioning this domain into simplices would require at least the $2^{d}$ function evaluations on the corners of the unit cube, and would therefore not scale well to large dimensions.
It is possible to get a certificate for integrals over $[0,1]^{d}$ by generalizing equation (11). The integrand there is nondecreasing and the input points are ‘biased low’ for the lower bound and ‘biased high’ for the upper bound. We can generalize the notion of points being biased low by requiring that $\delta({\bm{a}})\geqslant 0$ holds for all ${\bm{a}}\in[0,1]^{d}$. This property is known as ‘nonnegative local discrepancy’ (NNLD). An analogous ‘nonpositive local discrepancy’ (NPLD) property has $\delta({\bm{a}})\leqslant 0$ for all ${\bm{a}}\in[0,1]^{d}$.
The extension to general $d\geqslant 1$ requires a strong monotonicity condition. In particular, it is not enough for $f$ to be nondecreasing in each component of ${\bm{x}}$ individually with the other $d-1$ components held fixed. We require $f$ to be completely monotone as described by [1].
The function $f$ is completely monotone on $[0,1]^{d}$ if
$\sum_{v\subseteq u}(-1)^{|u-v|}f({\bm{x}}_{-v}{:}{\bm{z}}_{v})\geqslant 0$
holds for every nonempty $u\subseteq 1{:}d$ and all ${\bm{x}},{\bm{z}}\in[0,1]^{d}$ with ${\bm{x}}\leqslant{\bm{z}}$ componentwise. Taking $u=\{j\}$ this requires that $f({\bm{x}}_{-j}{:}{\bm{z}}_{j})\geqslant f({\bm{x}})$ so $f$ is nondecreasing in each of the $d$ variables. For $u=\{j,k\}$ for $j\neq k$ we see that $f({\bm{x}}_{-j}{:}{\bm{z}}_{j})-f({\bm{x}})$ is nondecreasing in $x_{k}$. Generally an $r$-fold difference of differences is nonnegative.
If we use ‘increasing’ as a less precise but more vivid term for ‘nondecreasing’, then we may say that $f$ is increasing in every variable, the amount by which it is increasing in any variable is increasing in any other variable, that amount in turn is increasing in any third variable, and so on. This is a strict requirement. It is satisfied if $f$ is the cumulative distribution function (CDF) of some random random vector ${\bm{X}}$, i.e., $f({\bm{x}})=\Pr({\bm{X}}\leqslant{\bm{x}}\ \text{componentwise})$. As a partial converse, if $f$ is also right continuous, then
$f({\bm{x}})=f({\bm{0}})+\lambda\nu([{\bm{0}},{\bm{x}}])$ (13)
holds for some probability measure $\nu$ and some $\lambda\geqslant 0$ [1].
Theorem 4.1.
Let $f$ be of the form in equation (13). If ${\bm{1}}-{\bm{x}}_{1},\dots,{\bm{1}}-{\bm{x}}_{n}$ have NNLD, then
$\frac{1}{n}\sum_{i=1}^{n}f({\bm{x}}_{i})\geqslant\int_{[0,1]^{d}}f({\bm{x}}){\mathrm{d}}{\bm{x}}.$
If ${\bm{x}}_{1},\dots,{\bm{x}}_{n}$ have NPLD and either all ${\bm{x}}_{i}\in[0,1)^{d}\cup\{{\bm{1}}\}$, or $\nu$ is absolutely continuous with respect to Lebesgue measure, then
$\frac{1}{n}\sum_{i=1}^{n}f({\bm{1}}-{\bm{x}}_{i})\leqslant\int_{[0,1]^{d}}f({\bm{x}}){\mathrm{d}}{\bm{x}}.$
Proof 4.2.
This is Theorem 1 of [39].
To get a certificate we can use $n$ points that have NNLD and $n$ points that have NPLD. Gabai [38] showed that the Hammersley sequence in $[0,1]^{2}$ (in any base $b\geq 2$) has NNLD. That is generalized in [39] to any digital net in $[0,1]^{d}$ where all of the generator matrices are permutation matrices. They also show that some rank one lattices have NNLD. The proofs of those NNLD results are based on the theory of associated random variables from reliability theory [35].
NPLD points are harder to construct than NNLD points. At first this is surprising. If $x_{1},\dots,x_{n}\in[0,1]$ have NNLD then they ‘oversample’ $[0,a)$ so they undersample $[a,1]$. Then $1-x_{i}$ undersample $[0,a]$ and therefore they undersample $[0,a)$. As a result, $1-x_{1},\dots,1-x_{n}\in[0,1]$ have NPLD. For $d\geqslant 2$ it is no longer true that NNLD points ${\bm{x}}_{i}$ must provide NPLD points ${\bm{1}}-{\bm{x}}_{i}$. The root of the problem is that extra points ${\bm{x}}_{i}\in[{\bm{0}},{\bm{a}})$ implies fewer points ${\bm{x}}_{i}\in[0,1]^{d}\setminus[{\bm{0}},{\bm{a}})$. But this set is not a hyperrectangle, and so neither is $\{{\bm{1}}-{\bm{x}}\mid{\bm{x}}\in[0,1]^{d}\setminus[{\bm{0}},{\bm{a}})\}$. The NPLD property requires undersampling of hyperrectangles containing the origin.
There are two basic constructions for NPLD points in [39]. For $d=1$, the points $x_{i}=i/n$ for $i=1,\dots,n$ have NPLD. For $d=2$ and Hammersley points ${\bm{x}}_{i}\in[0,1)^{2}$ the points $\tilde{\bm{x}}_{i}=({\bm{x}}_{i1}+1/n,1-{\bm{x}}_{i2})$ have NPLD. This ‘shift-flip’ transformation is from [26]. Tensor products of NNLD point sets have NNLD and tensor products of NPLD point sets have NPLD [39].
The greater challenge than finding NNLD and NPLD points is the requirement that $f$ be completely monotone. It can be weakened by finding a control variate function $g$ with $\int_{[0,1]^{d}}g({\bm{x}}){\mathrm{d}}{\bm{x}}=0$ such that $f+g$ is completely monotone, and then averaging $(f+g)({\bm{x}}_{i})$. This has not been explored in the literature. For some $f$ it would work to take $g({\bm{x}})=c\prod_{j=1}^{d}(x_{j}-1/2)$ for large enough $c>0$. Such a function $g$ is not ‘QMC friendly’ because its ANOVA decomposition has all of its variance in the highest-order $d$-dimensional term. As a result, for large $d$ we expect $(1/n)\sum_{i=1}^{n}g({\bm{x}}_{i})$ to converge slowly to $\mu$. For instance, $g$ has a very large Sobolev norm in the usual unanchored spaces. Using (3) for $\gamma_{j}=j^{-\eta}$ we get $\partial^{u}g=c\prod_{j\not\in u}(x_{j}-1/2)$ and then
$\|g\|_{{\bm{\gamma}}}^{2}$ $=c^{2}\sum_{u\subseteq 1{:}d}\frac{1}{\gamma_{u}}\int_{[0,1]^{|u|}}\biggl{(}\int_{[0,1]^{d-|u|}}\prod_{j\not\in u}(x_{j}-1/2){\mathrm{d}}{\bm{x}}_{-u}\biggr{)}^{2}{\mathrm{d}}{\bm{x}}_{u}$
$=c^{2}\sum_{u\subseteq 1{:}d}1_{u=1{:}d}\prod_{j\in u}j^{\eta}$
$=c^{2}(d!)^{\eta}.$
For $\eta=2+\varepsilon$ with $\varepsilon>0$, we get strong tractability. The number of function evaluations to get a QMC error for $g$ below $\epsilon$ can be brought below $c(d!)^{1+\varepsilon/2}$ times a constant that depends on $\epsilon$ but does not depend on $d$. The resulting rate in $d$ is quite unfavorable because $g$ has a very high norm in the weighted space.
Convergence rates for known NNLD and NPLD constructions show a dimension effect. The best constructions for even $d$ in [39] arise by taking tensor products of two dimensional Hammersley point sets on $m$ inputs (for NNLD points) along with tensor products of Hammersley point sets after the shift-flip transformation (for NPLD points). For $f\in\mathrm{BVHK}$ the resulting rule on $n=m^{d/2}$ points provides an error of $O(n^{-2/d+\epsilon})$. It is not known whether better convergence rates can be attained by other NNLD and NPLD constructions. For $d=2k+1$, one can either use a Cartesian product of $k+1$ two dimensional sets (ignoring one component) or use $k$ two dimensional sets and one endpoint rule.
While the $O(n^{-2/d+\epsilon})$ rate shows a strong dimension effect, it is better than the rate for deterministic sampling of bounded functions that are simply nondecreasing in each variable individually. Deterministic methods cannot be better than $O(n^{-1/d})$, and random methods cannot have a better RMSE than $O(n^{-1/d-1/2})$. See [87] for precise statements with proof.
The bracketing rules for convex functions that are twice continuously differentiable provide guaranteed intervals of width $O(n^{-2/d})$ for $\mu$. This is the optimal rate for integration of convex bounded functions. If =${\mathcal{F}}=\{f:[0,1]^{d}\to[0,1]\mid\text{$f$ is convex}\}$ then there exists $c_{d}>0$ such that any deterministic integration rule has $\sup_{f\in{\mathcal{F}}}|\hat{\mu}(f)-\mu(f)|\geqslant c_{d}n^{-2/d}$ when $n=(2m)^{d}/2$ for some integer $m\geqslant 1$ [56]. The deterministic rule are allowed to be adaptive, selecting ${\bm{x}}_{i}$ based on $f({\bm{x}}_{i^{\prime}})$ for $i^{\prime}<i$.
5 Randomized quasi-Monte Carlo
In RQMC our points ${\bm{x}}_{1},\dots,{\bm{x}}_{n}$ are individually ${\mathbb{U}}[0,1]^{d}$ while collectively having a small $D_{n}^{*}$. We can make $R$ statistically independent replicates $\hat{\mu}^{(1)},\dots,\hat{\mu}^{(R)}$ of $\mu$ using such an RQMC procedure. Then $\hat{\mu}_{n}^{(r)}=\hat{\mu}^{(r)}$ are IID with ${\mathbb{E}}(\hat{\mu}^{(r)})=\mu$. If the RQMC points have $D_{n}^{*}=O(n^{-1+\epsilon})$ then
$\hat{\mu}=\frac{1}{R}\sum_{r=1}^{R}\hat{\mu}^{(r)}$ (14)
has RMSE $O(R^{-1/2}n^{-1+\epsilon})$. If all of the mixed partial derivatives of $f$ taken at most once with respect to each of the $d$ variables are in $L_{2}[0,1]^{d}$, and we have scrambled some digital nets via the algorithm from [76] or from [69], then $\hat{\mu}$ from (14) has an RMSE that is $O(R^{-1/2}n^{-3/2+\epsilon})$. Even higher order rates in $n$ are available from randomizations of higher order digital nets [25].
These rates show that for a given number $nR$ of function evaluations, we expect a better estimate $\hat{\mu}$ by taking larger $n$ and smaller $R$. The RMSE of $\hat{\mu}$ describes the quality of our estimate of $\mu$ but not the quality of our UQ. Taking a larger value of $R$ will give better UQ, and this presents a fundamental tradeoff that we don’t have to consider in plain MC. Here we consider fixed $R$ as $n\to\infty$. Some comments on $R$ changing with $n$ are in Section 8.
We have two sample sizes to consider, $n$ and $R$. The customary sampling results in statistics have been developed for $n\to\infty$ independent observations. In RQMC we study $n\to\infty$, but for the very dependent observations used to create $\hat{\mu}_{n}^{(r)}$. Then there are $R$ replicates. While those are independent, we ordinarily prefer small $R$ instead of $R\to\infty$. The large sample size $n$ comes from dependent data and the independent sample size $R$ is small, so we do not have the large number of independent estimates that appears in most statistical theory.
We begin by presenting statistical results for $n\to\infty$ IID observations. Those results provides only a limited understanding of the accuracy of RQMC confidence intervals. That understanding combined with some knowledge of how RQMC works was used to design an extensive empirical investigation in [61]. Those empirical results have been followed by theoretical explanations of them and this is an active area of research.
Two important issues are how the error in the CLT (6) decreases with $n$ and how quickly the coverage of the standard Student’s $t$ confidence interval (9) approaches $1-\alpha$ as $n\to\infty$. These are most often studied through moment quantities. Let $\bar{Y}$ be the average of $n$ IID random variables $Y_{i}$ that have mean $\mu$, variance $\sigma^{2}>0$ and a finite third moment. Then by the Berry-Esseen theorem, there exists a constant $C<\infty$ such that
$\sup_{-\infty<z<\infty}\,\Bigl{|}\,\Pr\Bigl{(}\frac{\bar{Y}-\mu}{\sigma/\sqrt{n}}\leqslant z\Bigr{)}-\Phi(z)\Bigr{|}\leqslant\frac{C\rho}{\sigma^{3}\sqrt{n}}$
for $\rho=\mathbb{E}(|Y-\mu|^{3})$ and all $n\geqslant 1$. We can take $C=0.4748$ [91]. From this we see that the CLT takes hold at the $O(n^{-1/2})$ rate and that a scaled third central moment $\rho/\sigma^{3}$ governs the error. Usable confidence intervals, like the standard one, also have to contend with a generally unknown $\sigma$.
The error in confidence intervals like the standard one and also some bootstrap confidence intervals is commonly studied through scaled third and fourth moments
$\gamma=\frac{\mathbb{E}((Y-\mu)^{3})}{\sigma^{3}}\quad\text{and}\quad\kappa=\frac{\mathbb{E}((Y-\mu)^{4})}{\sigma^{4}}-3.$
These are known as the skewness and kurtosis of $Y$. A Gaussian random variable has $\gamma=\kappa=0$. If $\kappa<\infty$ and $Y_{i}$ are IID random variables with the same distribution as $Y$ then $\hat{\mu}=\bar{Y}=(1/n)\sum_{i=1}^{n}Y_{i}$ has skewness $\gamma/\sqrt{n}$ and kurtosis $\kappa/n$ [71]. Then the skewness of $\bar{Y}$ approaches 00 more slowly than the kurtosis does. A rapidly vanishing kurtosis is consistent with the third moment appearing in the Berry-Esseen bound but the fourth moment not appearing there.
The coverage error in the standard confidence interval for $\mu$ is
$\Pr\Bigl{(}\bar{Y}-\frac{s}{\sqrt{n}}t^{1-\alpha/2}_{(n-1)}\leqslant\mu\leqslant\bar{Y}+\frac{s}{\sqrt{n}}t^{1-\alpha/2}_{(n-1)}\Bigr{)}-(1-\alpha).$ (15)
This error has been well studied by researchers in the 1980s, especially Peter Hall [44, 45], in the context of bootstrap confidence intervals described below. A very interesting finding is that for IID sampling of $n$ observations, the coverage error in the standard interval is commonly $O(1/n)$ which is better than the RMSE of $O(n^{-1/2})$ for $\hat{\mu}$. The UQ achieves a better convergence rate than the estimate whose uncertainty it quantifies. The one-sided coverage errors $\Pr\bigl{(}\mu\leqslant\bar{Y}-{s}{t^{1-\alpha/2}_{(n-1)}/\sqrt{n}}\bigr{)}-\alpha/2$ and $\Pr\bigl{(}\mu\geqslant\bar{Y}+{s}t^{1-\alpha/2}_{(n-1)}/{\sqrt{n}}\bigr{)}-\alpha/2$ both converge at the $O(n^{-1/2})$ rate but a fortunate cancellation yields an $O(n^{-1})$ rate for (15).
For small values of $n$, some bootstrap confidence intervals can work better than the standard interval. The bootstrap is described in [34], [20] and [46] ranging from introductory to very technical. A bootstrap sample $Y_{1}^{*},\dots,Y_{n}^{*}$ is formed by taking $Y_{i}^{*}=Y_{j(i)}$ for $i=1,\dots,n$ where $j(i)\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}{\mathbb{U}}\{1,2,\dots,n\}$, so $Y_{i}^{*}$ are sampled with replacement from $Y_{1},\dots,Y_{n}$. We can then compute $\bar{Y}^{*}=(1/n)\sum_{i=1}^{n}Y_{i}^{*}$. That process can be repeated independently $B$ times ($B=1000$ is commonly used) yielding $\bar{Y}^{*1},\dots,\bar{Y}^{*B}$. We can sort those values getting $\bar{Y}^{*(1)}\leqslant\bar{Y}^{*(2)}\leqslant\cdots\leqslant\bar{Y}^{*(B)}$. The percentile confidence interval at the 95% level is $[\bar{Y}^{*(.025B)},\bar{Y}^{*(.975B)}]$. Nowhere does it use a parametric distributional assumption, such as the Gaussian distribution, for $Y_{i}$. It does require some moment assumptions in order to be asymptotically correct.
The bootstrap $t$ confidence interval, also called the percentile $t$ confidence interval has some advantages over the ordinary percentile interval. The error is typically $O(1/n)$ for both one-sided and two-sided intervals. For a distribution with $\gamma=\kappa=0$, the two-sided coverage error is $O(1/n^{2})$, so we expect it to work well for Gaussian or nearly Gaussian data. In some simulations [75] varying $\gamma$, $\kappa$ and $3\leqslant n\leqslant 20$, the bootstrap $t$ 95% confidence intervals had good coverage:
“The bootstrap $t$ method is shown to be very effective for the construction of central 95% confidence intervals for the mean of a small sample. Though it fails utterly when $n=3$, it achieves close to the nominal coverage for a diverse collection of continuous sampling distributions provided $n\geqslant 4$. The intervals can be quite long unless $n\geqslant 6$, and have a highly variable length unless $n\geqslant 7$.”
Of the 7 continuous distributions in [75], the bootstrap $t$ got close to nominal coverage in 6 of them. The exception was the lognormal distribution, where the bootstrap $t$ got only about 90% coverage by $n=20$. Even that was noticeably higher than all the other 8 methods apart from an alternative formulation of the bootstrap $t$.
The bootstrap $t$ method works as follows. From each bootstrap sample, we compute $t^{*}=\sqrt{n}(\bar{Y}^{*}-\bar{Y})/s^{*}$ where $(s^{*})^{2}=\sum_{i=1}^{n}(Y_{i}^{*}-\bar{Y}^{*})^{2}/(n-1)$. We do this $B$ times and then sort the resulting values getting $t^{*(1)}\leqslant t^{*(2)}\leqslant\cdots\leqslant t^{*(B)}$. The interval uses the approximation
$\Pr\Bigl{(}t^{*(0.025B)}\leqslant\sqrt{n}\frac{\bar{Y}-\mu}{s}\leqslant t^{*(0.975B)}\Bigr{)}\approx 0.95$
which leads to an approximate 95% confidence interval for $\mu$ of the form
$\Bigl{[}\bar{Y}-t^{*(0.975B)}\frac{s}{\sqrt{n}},\,\bar{Y}-t^{*(0.025B)}\frac{s}{\sqrt{n}}\Bigr{]}.$
Hall [45] gives some asymptotic expansions for the coverage error in confidence intervals for the mean. His Table 1 includes the percentile bootstrap, the bootstrap $t$ and the normal theory interval (8) that uses Gaussian quantiles instead of $t$ quantiles. Surprisingly, he does not include the standard Student’s $t$ intervals; for those see [80]. The coverage errors for two sided confidence intervals are
Normal theory: $\qquad(2/n)\varphi(z^{1-\alpha/2})\bigl{[}\phantom{-}0.14\kappa-2.12\gamma^{2}-3.35\bigr{]}+O(1/{n^{2}}),$
Student’s $t$: $\qquad(2/n)\varphi(z^{1-\alpha/2})\bigl{[}\phantom{-}0.14\kappa-2.12\gamma^{2}\,\,\,\phantom{-3.42}\bigr{]}+O(1/{n^{2}}),$
Percentile: $\qquad(2/n)\varphi(z^{1-\alpha/2})\bigl{[}-0.72\kappa-0.37\gamma^{2}-3.35\bigr{]}+O(1/{n^{2}}),\quad\text{and}$
Bootstrap $t$: $\qquad(2/n)\varphi(z^{1-\alpha/2})\bigl{[}-2.84\kappa+4.25\gamma^{2}\,\,\,\phantom{-3.42}\bigr{]}+O(1/{n^{2}}),$
where $\varphi$ is the probability density function of the standard Gaussian distribution. Hall’s names for the normal theory, percentile and bootstrap $t$ methods are, ‘Norm’, ‘BACK’ and ‘STUD’ respectively.
Hall’s assumptions include $8$ finite moments and a distribution for $Y_{i}$ whose support is not contained within an arithmetic sequence. His table is for nominal coverage $1-2\alpha$ whereas we ordinarily target coverage $1-\alpha$, but this only affects the scaling of the lead term, and not the qualitative consequences of skewness and kurtosis. Hall’s formulas for $n=18$ were very close to the average coverage for $16\leqslant n\leqslant 20$ in [75, Table 6], except for the lognormal distribution.
The coverage error formulas show an advantage for the bootstrap $t$. The term $4.25\gamma^{2}$ is positive, so it works to increase coverage. It also has no intercept, while the normal theory and percentile methods have a negative intercept which works to decrease their coverage. Simulations of coverage levels usually show that nonparametric confidence intervals have a coverage level that approaches the desired one from below as $n\to\infty$. That is, undercoverage is more common than overcoverage. Student’s $t$ intervals have a positive coefficient for kurtosis, a negative coefficient for squared skewness and no intercept.
The distribution of the $t$ statistic under non-normality is much studied. For IID data from a distribution symmetric about $\mu$ with heavier tails than the Gaussian, there is a tendency for confidence intervals based on the $t$ statistic to be conservative, though precise statements of this phenomenon require some extra steps. An explanation and survey of this work, going back nearly 100 years, appears in [17]. At a high level, higher kurtosis in $Y_{i}$ brings a lower kurtosis in $t$, so that extreme thresholds are exceeded by $|t|$ less often. What happens is that very large values of $|Y_{i}-\mu|$ inflate the denominator $s/\sqrt{n}$ more than the numerator $\bar{Y}-\mu$. In the extreme, if we send $Y_{1}\to\pm\infty$ then $t\to\pm 1$. The consequence is that for distributions with large kurtosis, the usual 95% confidence intervals, based on $|t|\leqslant t_{(n-1)}^{0.975}$ can cover the mean more often than the nominal level, because $t_{(n-1)}^{0.975}\geqslant 1.96$.
To switch from folklore to precise statements requires some additional care and caveats. Here are a few of those results. One subtlety is that the tendency to overcoverage holds at customary confidence levels, but not necessarily at lower levels, such as those below 50%. The distribution of $t$ does not depend on $\mu$ so it is studied with $\mu=0$. Then the findings are mostly studied through $\tilde{t}=\tilde{t}_{n}=\sum_{i=1}^{n}Y_{i}/(\sum_{i=1}^{n}Y_{i}^{2})^{1/2}$, with $t=\tilde{t}(n-1)^{1/2}/(n-\tilde{t}^{2})^{1/2}$, a monotone transformation. For even $\nu\geqslant 2$, ${\mathbb{E}}(\tilde{t}^{\nu})\leqslant{\mathbb{E}}(Z^{\nu})$ for $Z\sim\mathcal{N}(0,1)$; see Corollary 1 of [32]. Corollary 2 shows that $\tilde{t}$ has negative kurtosis under conditions that include IID symmetrically distributed $Y_{i}$. If the random variables are distributed as a scale mixture of mean zero Gaussians (i.e., $\mathcal{N}(0,\sigma^{2})$ for random $\sigma$) then the $t$ test is conservative at thresholds above $1.8$ [7]. More precisely: the bound was established for $2\leqslant n\leqslant 18$, with a critical threshold that never went above $1.8$ and decreased with increasing $n>6$. Most of the results are for symmetrically distributed $Y_{i}$, but it is known that the skewness of $t$ is $-2\gamma/\sqrt{n}+O(n^{-3/2})$ where $\gamma$ is the skewness of $Y_{i}$ [71, Equation (1.15)]. For a survey of related results, see [90].
Now we switch to the RQMC context. We replace $n$ by our number $R$ of replicates and $Y_{i}$ by $\hat{\mu}_{n}^{(r)}$. Only a little is known about the skewness and kurtosis of $\hat{\mu}_{n}$ for various RQMC methods and much of that knowledge is recent. Let these be $\gamma_{n}$ and $\kappa_{n}$, respectively.
First, there is an old result [67] proving a CLT as $n\to\infty$ for the scrambled $(t,m,d)$-nets using the scrambling from [76] when the underlying digital net has $t=0$, under smoothness conditions on $f$. That setting then has $\gamma_{n}\to 0$ and $\kappa_{n}\to 0$ and we expect the standard intervals should then have low coverage error. The nets of Faure [36] have $t=0$ but the more widely used ones based on sequences from Sobol’ [94] only have $t=0$ for $d\leqslant 2$ [74, Table 4.1]. On the other hand, randomly shifted lattice rules are known to produce $\hat{\mu}$ that does not follow a CLT [60]. There the distribution of $\hat{\mu}$ has a density represented by a spline curve.
In the present context with $R$ replicates, we expect the coverage error to be $O(1/R)$ as $R\to\infty$ for fixed $n$. Those results don’t tell us about what happens when $n\to\infty$ for fixed $R$.
The RQMC counterpart to our sample variance $s^{2}$ of (7) is
$\tilde{s}^{2}=\tilde{s}_{n}^{2}=\frac{1}{R-1}\sum_{r=1}^{R}(\hat{\mu}^{(r)}-\hat{\mu})^{2}$ (16)
and then the standard interval becomes $\hat{\mu}\pm\tilde{s}t^{1-\alpha/2}_{(R-1)}.$ This interval would be exact if $\hat{\mu}^{(r)}$ had a Gaussian distribution, and we can therefore expect it to work well when a CLT applies to $\hat{\mu}_{n}^{(r)}$ as $n\to\infty$.
The simulations in [61] investigated three different confidence interval methods for RQMC: the standard interval, the percentile bootstrap and the bootstrap $t$. The number of replicates was $R\in\{5,10,20,30\}$ in keeping with the desire to keep $R$ small. There were 5 different RQMC methods. Two of them were lattice rules from LatNet Builder [59] using random shifts modulo one, with and without the baker transformation of [48]. The other three RQMC methods used Sobol’s digital nets [94] using the direction numbers from [55]. They were randomized with either a digital shift (see [58]), a matrix scramble of [69] plus digital shift, or the nested uniform scramble from [76]. The sample sizes were $n=2^{m}$ for $m\in\{6,8,10,12,14\}$. There were $6$ integrands with known $\mu$ chosen to mix cases that are easy and difficult for RQMC, varying in their levels of smoothness and varying in the extent to which they have low effective dimension. See [61] for a discussion. An integrand that is easy to integrate well is not necessarily one that makes the confidence interval problem easy. All of those integrands were constructed to allow varying dimension $d$. The simulation used $d\in\{4,8,16,32\}$.
Each confidence interval method was challenged with $2400$ use cases from $6$ integrands, $4$ dimensions, $5$ sample sizes, $5$ RQMC methods and $4$ values of $R$. Each challenge was repeated $1000$ times by taking $R$ sample values without replacement from a pool of $10{,}000$. The goal was to get 95% coverage. A method was deemed to fail if it would only attain 94% coverage. If it covers $\mu$ less than 927 times out of 1000, that would happen with less than 0.04 probability (by the binomial distribution) for a method that had coverage 0.94 or more. As a result any such setting was deemed to be a confirmed failure of the RQMC confidence interval.
Of the 2400 cases, the percentile bootstrap was found to fail 1689 times. This is not surprising given the results in [75]. The bootstrap $t$ was found to fail 81 times. The standard interval failed only 3 times. None of those failures were for $R=10$. Figure 2 of [61] erroneosly shows one such failure. That happened because the plotting command to set up the figure’s axes mistakenly did not use type =“n” in the call to the plot function in the R language.
Based on the results in [61], the current best recommendation for 95% RQMC confidence intervals is to use $R\geqslant 10$ independent replications along with the standard Student’s $t$ based confidence interval, along with one’s preferred RQMC method. Next we relate this finding to the understanding of confidence intervals from Hall’s formulas.
From the data in [61] it was possible to compute sample values of the skewness and kurtosis in each distribution of $\hat{\mu}$ based on $10{,}000$ evaluations for each integrand, dimension $d$, sample size $n$ and RQMC method. Inspection of those estimated values revealed that most of the RQMC distributions had modest sample skewnesses and many of them had very large sample kurtoses. A large kurtosis presents a difficulty for the bootstrap $t$ and an advantage for the standard interval as discussed above. The modest skewness removes a disadvantage for the standard interval and it removes an advantage for the bootstrap $t$. Figure 2 of [61] shows that in some instances with extremely large kurtosis, the standard confidence interval had coverage well above the nominal 95% level. This consequence of high kurtosis is well known [17] as mentioned above.
The empirical findings motivated subsequent work in [85] which shows that for random linear scrambling of a digital net in base $2$ with a random digital shift, the skewness of $\hat{\mu}_{n}$ is $\gamma_{n}=O(n^{\epsilon})$ for any $\epsilon>0$, so it is almost $O(1)$. Furthermore, under a model with randomly chosen generator matrices the skewness is $O(n^{-1/2+\epsilon})$. The kurtosis of $\hat{\mu}$ from random linear scrambling of a digital net in base $2$ is known to diverge to $\infty$ as $n\to\infty$ for an analytic function on $[0,1]$ by a finding in [84, Section 3]. The reasoning is as follows. There is an event of probability $\Omega(1/n)$ that gives a squared error of $\Omega(1/n^{2})$ (for smooth $f^{\prime}$ that does not integrate to zero). That same event on its own contributes $\Omega(n^{-5})$ to the fourth power of the error. The expected squared error is $O(n^{-3+\epsilon})$ and so $\kappa_{n}+3=\Omega(n^{-5})/O(n^{-6+2\epsilon})=\Omega(n^{1-2\epsilon})$. An integrand on $[0,1]^{d}$ for $d>1$ with a smooth contribution from a one dimensional main effect will have the same diverging kurtosis and the critical event has probability $\Omega(d/n^{2})$. Having $\kappa_{n}\to\infty$ completely rules out a CLT for $\hat{\mu}_{n}$ from matrix scrambling with a digital shift.
The empirical findings of modest skewness and potentially very large kurtosis have not been established for the other RQMC methods. The other methods in [61] did not appear to have large skewness. Only three of 600 cases had $|\hat{\gamma}|>4$. Those had large kurtoses and since the variance of $\hat{\gamma}$ involves sixth moments, they might just be sampling fluctuations. Many of the histograms of $\hat{\mu}$ had non-Gaussian but nearly symmetric distributions.
The bias-corrected accelerated bootstrap (BCa) of [33] was not included in the simulations of either [61] or [75]. Hall [45] prefers the bootstrap $t$, while DiCiccio and Efron [23] prefer the BCa. Table 1 of [80] includes an entry ‘ABC’ for a method that is very close to BCa. The BCa has a coverage expression of $-2.68\kappa+3.17\gamma^{2}-3.42$ (after corrections). The constant $-3.42$ and the $-2.68\kappa$ term both suggest that BCa is not well suited to the RQMC estimates that can have very large kurtosis and minimal skewness.
6 The Warnock-Halton quasi-standard error
Tony Warnock and John Halton reasoned that QMC ideas should also be usable to get not just an estimate of $\mu$ but also an estimate of the uncertainty in an estimate of $\mu$. Instead of using $R$ completely random estimates (by Monte Carlo) they use QMC ideas to balance those $R$ replicates with the goal of getting ‘better than random’ replication. Their proposal is in the technical reports [99, 100] and article [47].
To get $R\geqslant 2$ quasi-replicates of an integral on $[0,1]^{d}$, Warnock [99] takes QMC points ${\bm{x}}_{i}\in[0,1]^{dR}$. Then for $r=1,\dots,R$ let
$\tilde{\bm{x}}_{i,r}=(x_{i,d(r-1)+1},x_{i,d(r-1)+2},\cdots,x_{i,dr})\in[0,1]^{d}$
so ${\bm{x}}_{i}=(\tilde{\bm{x}}_{i,1},\tilde{\bm{x}}_{i,2},\cdots,\tilde{\bm{x}}_{i,R})$. Now let
$\hat{\mu}^{(r)}=\frac{1}{n}\sum_{i=1}^{n}f(\tilde{\bm{x}}_{i,r})$
for $r=1,\dots,R$. Then $\hat{\mu}=(1/R)\sum_{r=1}^{R}\hat{\mu}^{(r)}$ and its quasi-standard error is
$\text{QSE}=\biggl{(}\frac{1}{R}\frac{1}{R-1}\sum_{r=1}^{R}(\hat{\mu}^{(r)}-\hat{\mu})^{2}\biggr{)}^{1/2}.$
Halton [47] reasons that when ${\bm{x}}_{1},\dots,{\bm{x}}_{n}$ have very low discrepancy then $\tilde{\bm{x}}_{i,r}$ is effectively independent of $\tilde{\bm{x}}_{i,r^{\prime}}$ for $1\leqslant r<r^{\prime}\leqslant R$. That holds for random $i\sim{\mathbb{U}}\{1,\dots,n\}$ with any two values $r$ and $r^{\prime}$ held fixed because ${\bm{x}}_{1},\dots,{\bm{x}}_{n}$ have nearly the ${\mathbb{U}}[0,1]^{dR}$ distribution. Then $f({\bm{x}}_{i,r})$ and $f({\bm{x}}_{i,r^{\prime}})$ are effectively independent of each other too. This does not however make $\hat{\mu}^{(r)}$ and $\hat{\mu}^{(r^{\prime})}$ effectively independent. For that, it would suffice to have $(\tilde{\bm{x}}_{1,r},\dots,\tilde{\bm{x}}_{n,r})$ effectively independent of $(\tilde{\bm{x}}_{1,r^{\prime}},\dots,\tilde{\bm{x}}_{n,r^{\prime}})$, but small $D_{n}^{*}$ does not imply this.
The QSE can be far too small [78]. For instance if ${\bm{x}}_{i}$ are points of a Sobol’ sequence and $f$ is additive, then we will get a QSE of zero. That happens because $\{x_{1j},x_{2j},\dots,x_{nj}\}=\{0,1/n,\dots,(n-1)/n\}$ for all $j=1,\dots,dR$ and then $\hat{\mu}^{(r)}=\hat{\mu}^{(1)}$ for $r=2,\dots,R$. Some good results forming confidence intervals based on the QSE are reported in [100]. That source may be hard to find; the results are described in [78].
Ideas like the QSE have potential to bring QMC accuracy with respect to $R\to\infty$ for uncertainty quantification based on $R$ replicates instead of the customary MC rate in $R$. The near independence that Halton writes about could perhaps be achieved using a QMC point set $\tilde{\bm{X}}\in{\mathbb{R}}^{R\times nd}$. We can rearrange row $r$ of $\tilde{\bm{X}}$ into an $n\times d$ matrix to use for the $r$’th QMC replicate. Finding a good QMC point set with $R$ rows and $nd$ columns presents a substantial challenge for the values of $n$, $R$ and $d$ commonly used in RQMC. There may however be some alternative way to get QMC accuracy with respect to $R$ using something more complicated than their procedure that is also less cumbersome than finding $R$ QMC points in dimension $nd$. Their proposal has not been explored or modified very much.
7 Guaranteed automatic integration library
Most of the uncertainty quantification methods in this article are about finding some value $\epsilon$ where, after sampling $f$ $n$ times, we have reason to believe that $|\hat{\mu}_{n}-\mu|\leqslant\epsilon$. The evidence behind that belief might be a certificate, a confidence interval or an asymptotic confidence interval. A complementary approach is to start with a target value of $\epsilon$ and look for a value of $n$ for which we will have reason to believe that $|\hat{\mu}_{n}-\mu|\leqslant\epsilon$. There are many results in the literature showing how $n$ must grow asymptotically as $\epsilon$ is reduced. The Guaranteed Automatic Integration Library (GAIL) [96] may be the unique one that provides a non-asymptotic solution for integration over $[0,1]^{d}$. The GAIL project continues as part of the qmcpy Python library [15].
In GAIL, the user specifies $\epsilon$ and then the library is designed to return an interval $[a,b]$ of width no more than $2\epsilon$ with $a\leqslant\mu\leqslant b$. In other words, GAIL provides a bracketing solution where the user can specify the width of the bracket instead of specifying $n$. A probabilistic version delivers an interval with $\Pr(a\leqslant\mu\leqslant b)\geqslant 1-\alpha$.
A first approach to GAIL using MC is in [50]. As noted in Section 3 such an approach requires some extra knowledge about $f({\bm{x}})$. The approach in [50] is to use a preliminary sample to get a probabilistic upper bound for $\sigma$. If $\Pr(\sigma\leqslant\hat{\sigma})\geqslant 1-\alpha_{1}$ for $\alpha_{1}<\alpha$ then we only need to find $a<b$ with $\Pr(a\leqslant{\mathbb{E}}(f({\bm{x}}))\leqslant b)\geqslant 1-\alpha_{2}$ that holds whenever $\operatorname{Var}(f({\bm{x}}))\leqslant\hat{\sigma}^{2}$ and $\alpha_{1}+\alpha_{2}\leqslant\alpha$.
Now we need an upper confidence limit for $\sigma$. This seems like it should be even harder to get than the confidence interval for $\mu$ that we set out to get and it also requires extra information. The extra information used in [50] is the assumption that
${\mathbb{E}}((f({\bm{x}})-\mu)^{4})\leqslant\tilde{\kappa}\sigma^{4}$ (17)
for known $\tilde{\kappa}<\infty$. That assumption allows for a first Monte Carlo sample with $n_{1}\geqslant 2$ function evaluations to give an upper confidence bound $\hat{\sigma}^{2}$ for $\sigma^{2}$. That is followed by a second independent Monte Carlo sample to provides a confidence interval for $\mu$. The second sample size $n_{2}$ can now be chosen using $\hat{\sigma}^{2}$ to get the desired interval length. The first stage uses a probability inequality of Cantelli and the second uses a Berry-Esseen inequality where the third absolute moment $\rho$ is bounded using $\tilde{\kappa}$.
The assumption (17) is discussed in [50]. It corresponds to a cone of integrands
$\bigl{\{}f\in L_{4}[0,1]^{d}\mid\|f-\mu(f)\|_{4}\leqslant\tilde{\kappa}^{1/4}\|f-\mu(f)\|_{2}\bigr{\}}.$ (18)
Scaling $f$ by a constant factor keeps it within the cone. That is quite different from using a ball of functions instead, because scaling a function can move it in or out of a ball. The cone (18) is non-convex, so it avoids condition (iii) in the impossibility result of Bahadur and Savage. The cost of the algorithm is not very sensitive to overestimation of $\tilde{\kappa}$. If $f({\bm{x}})$ has a Gaussian distribution with $\sigma>0$ then $\tilde{\kappa}=3$. No distribution has $\tilde{\kappa}<1$.
A QMC version of GAIL is in [51] . It is based on the Sobol’ sequence and a Walsh decomposition of the integrand. See [28] for both of those. The Walsh decomposition expands $f$ into a sum of Walsh functions indexed by ${\bm{k}}\in\{0,1,2,\dots\}^{d}$ and multiplied by coefficients $\hat{f}_{{\bm{k}}}$. The integration error can be bounded by a sum of $|\hat{f}_{{\bm{k}}}|$ over certain indices ${\bm{k}}$ that depend on which Sobol’ points are used. Given a sample from a Sobol’ sequence it is possible to estimate some of the Walsh function coefficients. They make an assumption about how Walsh coefficients decay as ${\bm{k}}$ moves farther from $\bm{0}$. Their assumption goes beyond the decay of Walsh coefficients noted by [24] and [102] for smooth $f$. That assumption places $f$ inside a cone and it lets them bound the sum of the remaining absolute coefficients given the estimated sum of absolute coefficients. They keep doubling the number of Sobol’ points and updating the error bound until the error bound is small enough.
A counterpart for rank one lattices, using a Fourier decomposition is in [54]. A discussion of relative error is studied in [95].
8 New directions
We cannot always write an expectation $\mu$ of interest as an integral of some computable function $f$ over $[0,1]^{d}$ for finite $d$, even with all the methods of [21] at our disposal. For instance, many problems in Bayesian computation have $\mu={\mathbb{E}}(g({\bm{x}}))$ for ${\bm{x}}\sim p$ where we cannot readily sample independently from the posterior distribution $p$. This $p$ may depend on arbitrary details of a very large data set. This sampling difficulty is the main impetus for methods like Markov chain Monte Carlo (MCMC) [10] and particle filters [16]. Because MCMC and particle methods sample dependently it is more difficult to quantify the uncertainty in their estimates.
Unbiased MCQMC
A second challenge with UQ for MCMC is that their estimates typically have a bias. That bias may disappear exponentially fast with $n$ while remaining large in practice because $A\rho^{n}$ for $\rho$ just barely smaller than one may be large for the $n$ we can use. It is much easier to quantify uncertainty in unbiased estimates that can be independently replicated. Coupling from the past (CFTP) [88] can generate unbiased estimates in some MCMC settings, but it only works well in very restrictive settings. There is some recent work on more generally applicable coupling strategies that remove the bias from MCMC. See [3] for a discussion of those methods. That allows independent replicates of unbiased MCMC methods to be used for UQ.
It is possible to embed QMC and RQMC methods into MCMC by replacing the sequence of IID random numbers driving the MCMC sampling by a sequence that is completely uniformly distributed (CUD) [14, 11, 97, 82, 12, 65]. Using a CUD sequence is like using a small pseudo-random number generator in its entirety. CUD sequences are described in [64] and a probabilistic version called weakly CUD is given in [98]. These methods are proven to converge to $\mu$ under assumptions similar to those where usual MCMC converges. Faster convergence is usually observed especially for the Gibbs sampler. Chen [11] establishes an RMSE of $o(n^{-1/2})$, under strong assumptions. These methods can be replicated but as for IID MCMC sampling, replication does not let us eliminate the effects of bias.
By combining unbiased MCMC with a weakly CUD driving sequence, [31] are able to get independent unbiased replicates of MCMC estimates for the Gibbs sampler. That allows simple replication for MCMC with empirically better convergence than by plain MCMC. Much of the analysis extends beyond the Gibbs sampler, but the Gibbs sampler is smooth which can help it benefit from more evenly distributed inputs and their algorithm also uses a coupling strategy designed for the Gibbs sampler. When CFTP is applicable, it can be used with RQMC [62].
Normalizing flows
There has been much recent interest in normalizing flows [89] that produce a transformation $\phi$ of ${\bm{x}}\sim{\mathbb{U}}[0,1]^{d}$ so that ${\bm{z}}=\phi({\bm{x}})\sim q$ where $q\approx p$. Usually $\phi$ is a transformation of $d$ Gaussian random variables, but those can be expressed as a transformation of $d$ uniform ones. Then we may estimate $\mu$ by a self-normalized importance sampling estimate
$\hat{\mu}=\frac{1}{n}\sum_{i=1}^{n}g({\bm{z}}_{i})\frac{p({\bm{z}}_{i})}{q({\bm{z}}_{i})}\Bigm{/}\frac{1}{n}\sum_{i=1}^{n}\frac{p({\bm{z}}_{i})}{q({\bm{z}}_{i})}$
where ${\bm{z}}_{i}=\phi({\bm{x}}_{i})$ for ${\bm{x}}_{i}\sim{\mathbb{U}}[0,1]^{d}$. We need to be able to compute $p$ up to a normalizing constant. We need to sample from $q$ and also evaluate an unnormalized version of it, but we are free to choose $q$ from a flexible parametric family for which these are feasible. It is also required that $q({\bm{z}})>0$ whenever $p({\bm{z}})>0$.
The use of normalizing flows allows MC and RQMC methods to compute confidence intervals unaffected by the bias in MCMC. The first effort using RQMC in normalizing flows is in [2]. The results there show decreasing effectiveness as the dimension increases. It is reasonable to suppose that the ratio $p/q$ becomes very unfavorable to RQMC as the dimension increases. Some more favorable results are in [66] which uses transformations tuned to RQMC along with some dimension reduction ideas.
Median of means
The distribution of RQMC estimates does not generally follow a CLT. When the matrix scramble of [69] with a digital shift is applied to a Sobol’ net, the distribution of $\hat{\mu}_{n}-\mu$ can be very non-Gaussian for a smooth integrand. As noted above, the kurtosis of $\hat{\mu}_{n}$ may diverge to infinity as $n$ increases, while the skewness remains modest or even converges to zero. In that setting, much of the variance is due to rare events of probability $\Omega(1/n)$ where $|\hat{\mu}-\mu|=\Omega(n^{-1})$. If one takes the median of $R$ replicates instead of the mean, then the outliers among $\hat{\mu}^{(1)},\dots,\hat{\mu}^{(R)}$ are essentially ignored. This is called a ‘median of means’ because each $\hat{\mu}^{(r)}$ is the mean of $n$ function evaluations, and we then take $\hat{\mu}=\mathrm{median}(\hat{\mu}^{(1)},\dots,\hat{\mu}^{(R)})$.
For analytic functions on $[0,1]^{d}$, a median-of-means approach brings an error of $O(n^{-c\log_{2}(n)/d})$ for any $c<3\log(2)/\pi^{2}\approx 0.21$ [86]. This rate is called super-polynomial because it is better than $O(n^{-r})$ for any finite $r$. A median-of-means strategy adapts to a possibly unknown level of smoothness in $f$. For $R\geqslant\log_{2}(n)$, it attains an RMSE of $O(n^{-\alpha-1/2+\epsilon})$ for any $\epsilon>0$ when $f$ has finite variation of order $\alpha$ (defined in [25]). This is proved in Theorem 2 of [83].
There are some other applications of the median of means method that attain a universal goodness property. Here are sketches of two of them; the reader should read them to get the full details. A median of means approach to RQMC by lattice rules has been developed by [42]. Lattice rules can be tuned to specific Hilbert space weights. The median rules in [42] can be constructed without specifying those weights and yet they attain nearly the optimal worst-case error rate for any reasonable choice of weights and smoothness. Goda and Krieg [41] choose a random prime number $p$ from the interval $[\lfloor n/2\rfloor+1,n]$. Then they choose a rank one lattice in $[0,1]^{d}$ with $p$ points using a generating vector chosen uniformly from $\{1,2,\dots,p-1\}^{d}$. From that rank one lattice, they estimate $\mu$. They repeat this $R=\Omega(\log(n))$ times independently and take $\hat{\mu}=\mathrm{median}(\hat{\mu}^{(1)},\dots,\hat{\mu}^{(R)})$. The resulting $\hat{\mu}$ is nearly optimal in any Korobov class of periodic functions with smoothness $\alpha>1/2$.
The great promise of median-of-means estimates strongly motivates us to seek confidence intervals or some other uncertainty quantification for them. Suppose that $\hat{\mu}_{r}$ has a continuous distribution with a unique median $\tilde{\mu}$. Then $\Pr(\hat{\mu}_{r}<\tilde{\mu})=\Pr(\hat{\mu}_{r}>\tilde{\mu})=1/2$. It is straightforward to sort the $\hat{\mu}^{(r)}$ into $\hat{\mu}^{[1]}\leqslant\hat{\mu}^{[2]}\leqslant\cdots\leqslant\hat{\mu}^{[R]}$ and from that get a confidence interval on $\tilde{\mu}$. For instance $\Pr(\tilde{\mu}<\hat{\mu}^{[r]})=\sum_{\ell=0}^{r-1}{R\choose\ell}/2^{R}$ can be used to get an upper confidence limit for $\tilde{\mu}$ (depending on $r$) and a lower limit can be attained similarly. This falls short of providing a confidence interval for $\mu$ because there is no assurance that $\tilde{\mu}=\mu$. We would need a computable bound (probabilistic or otherwise) for $|\tilde{\mu}-\mu|$ in order to get a non-asymptotic uncertainty quantification for $\mu$ this way.
The median of means method is most commonly used on $n$ IID random variables $Y_{i}$ with finite variance $\sigma^{2}$. The emphasis is very different from our use taking the median of $R$ independent estimates that each use $n$ very dependent values. While a CLT gives asymptotic confidence intervals for $\mu={\mathbb{E}}(Y_{i})$ the goal in much median of means research is to get the desired coverage for finite $n$. That requires somewhat wider confidence interval than the CLT based ones and also is only available for $\alpha\geqslant\alpha_{\min}>0$ for a threshold $\alpha_{\min}$ that depends on some assumptions about the distribution of $Y_{i}$. See [22] and references therein. The methods use known $\sigma$ though an upper bound or upper confidence limit for $\sigma$ (as in GAIL) could be used.
We could split our $R$ replicates into a small number $b$ of subsets of $R/b$ replicates, and take the median of $b$ subset means. We could then get a finite $R$ confidence interval for $\mu$, but the ones in [22] require knowledge of $\sigma^{2}_{n}=\operatorname{Var}(\hat{\mu}_{n}^{(r)})$ or an upper bound for that variance. They also have width proportional to $\sigma_{n}$ while the median of means estimates in [86, 84, 83] have error $o(\sigma_{n})$. Therefore a median of means confidence interval would be quite conservative. Gobet et al. [40] compare several robust confidence interval strategies for RQMC, including median of means, for known $\sigma_{n}$, but do not declare a winning method.
$R$ growing with $n$
For $N=Rn$ function evaluations, we could take $n=N^{c}$ and $R=N^{1-c}$ (i.e., integer values near these) for $10<c<$ 1 as $N\to\infty$ as studied in [72]. To remove an uninteresting complication, they assume that $\sigma_{n}>0$ for all $n$. They develop CLTs for $\hat{\mu}=(1/R)\sum_{r=1}^{R}\hat{\mu}_{N/R}^{(r)}$. A CLT along with an estimate $\hat{\sigma}^{2}_{n}$, such as $\tilde{s}_{n}^{2}$ of (16), which satisfies $\lim_{n\to\infty}\Pr\bigl{(}|\hat{\sigma}_{n}^{2}/\sigma^{2}_{n}-1|>\epsilon\bigr{)}=0$ for any $\epsilon>0$ provides an asymptotically valid confidence interval for $\mu$ via (8) or (9) (with $R$ playing the role of the sample size in those equations).
Even though $\hat{\mu}_{N/R}^{(r)}$ are IID for fixed $N$, that common distribution changes as $N\to\infty$ and so the CLT has to be of the triangular array type. That requires a Lindeberg condition (see page 13:10 of [72]) for which a simpler Lyapunov condition is sufficient. The Lyapunov condition is that
$\frac{{\mathbb{E}}\bigl{(}|\hat{\mu}_{N/R}-\mu|^{2+\delta}\bigr{)}}{R^{\delta/2}\sigma^{2+\delta}_{N/R}}=\frac{{\mathbb{E}}\bigl{(}|\hat{\mu}_{N^{c}}-\mu|^{2+\delta}\bigr{)}}{N^{(1-c)\delta/2}\sigma^{2+\delta}_{N^{c}}}\to 0$ (19)
as $N\to\infty$, for some $\delta>0$. A sufficient condition for (19) is that ${\mathbb{E}}(|\hat{\mu}_{n}-\mu|^{2+\delta})/\sigma_{n}^{2+\delta}<k$ holds for some $k$ and all sufficiently large $n$. Table 2 of [72] gives upper bounds on $c$ to get a CLT under various assumptions on the regularity of $f$ and the Lyapunov conditions. There are also some upper bounds under the additional requirement that $\hat{\sigma}_{n}/\sigma_{n}$ converges in probability to one.
For matrix scrambling with a digital shift and a smooth enough integrand, we have ${\mathbb{E}}(|\hat{\mu}_{n}-\mu|^{2+\delta})=\Omega(n^{-3-\delta})$ by the argument in Section 5 and $\sigma^{2}_{n}=O(n^{-3+\epsilon})$. Then
$\frac{\Omega(n^{-3-\delta})}{R^{\delta/2}O(n^{(-3+\epsilon)(1+\delta/2)})}=\Omega\Bigl{(}\frac{n^{\delta/2-\epsilon(1+\delta/2)}}{R^{\delta/2}}\Bigr{)}$
for any $\epsilon>0$. As a result, (19) cannot hold for any $\delta>0$ if $R=o(n)$.
Other RQMC methods and other smoothness conditions can impose less stringent requirements on $R$ for a CLT. For example the digital nets of [36] scrambled as in [76] are known to yield a CLT for $\hat{\mu}$ as $n\to\infty$ for fixed finite $R$ [67]. When UQ is the primary criterion with accuracy secondary, then those $t=0$ nets might be preferable to the ones of Sobol’ with $t>0$ where no CLT has been proven.
9 Conclusions
We have seen that uncertainty quantification for QMC estimates is subject to a complex mix of gaps and tradeoffs and surprises. The bracketing inequalities for smooth convex integrands are simple to use and very easy to understand. However they require strong assumptions, exhibit a severe dimension effect and can greatly overestimate the size of the error. The other methods with certificates also require strong assumptions and show a dimension effect. The methods based on NNLD and NPLD points use more involved constructions and more complex derivations.
We may be able to get a certificate that becomes narrow at the customary convergence rate if we know $V_{\mathrm{HK}}(f)$ or a weighted Hilbert space norm for $f$. However such knowledge is likely to be extremely rare. An unanchored weighted Hilbert space norm for $f$ also includes a term $\bigl{(}\int_{[0,1]^{d}}|\partial^{1:d}f({\bm{x}})|^{2}{\mathrm{d}}{\bm{x}}/\gamma_{1:d}\bigr{)}^{1/2}$ which can be large enough to make such an error estimate very conservative.
When we are willing to accept a confidence interval instead of a certificate, then more methods become available. These require strong assumptions such as a known bound on $f$ or a known cone to which $f$ must belong. Confidence intervals that are still correct for any $f$ in a large class of integrands can be very conservatively wide.
If we are willing to accept an asymptotic confidence interval, then more choices become available. A plain Student’s $t$ confidence interval based on a modest number $R$ of random replicates of an RQMC rule performed well in simulations. One explanation, which needs more study, is that $\hat{\mu}_{n}$ from at least some RQMC methods takes on a more symmetric distribution as $n\to\infty$.
The phenomenon of $\kappa_{n}\to\infty$ poses extreme difficulty for $\tilde{s}^{2}_{n}$ to converge to $\sigma^{2}_{n}$, which is one of the conditions that [72] use for asymptotic confidence intervals. That would require $\kappa_{n}/R\to 0$, but we want small $R$ for accurate estimation of $\mu$. When $\kappa_{n}\to\infty$ due to rare outliers, then we may not see any of those outliers among $R$ replicates. We will then get an estimate $\tilde{s}^{2}$ that is far smaller than $\sigma^{2}_{n}$. Ordinarily that would be very unfavorable for confidence interval coverage. However if the distribution of $\hat{\mu}$ is symmetric or close enough to symmetric, then the standard interval will give reliable and even somewhat conservative coverage despite the frequent underestimation of $\sigma_{n}$.
It is a pleasant surprise that some RQMC estimates tend towards symmetry for large $n$ making it possible to get reliable confidence intervals for $\mu$ without having a CLT or a good estimate of $\sigma^{2}_{n}$. This clearly needs more study, to see what conditions we need on the integrands and RQMC methods for this to happen.
Acknowledgments
This work was supported by the National Science Foundation under grant DMS-2152780. I thank the people behind MCQMC 2024: Christiane Lemieux, Ben Feng, Nathan Kirk, Adam Kolkiewicz, Carla Daniels and Greg Preston for organizing such a delightful meeting. Most of the new work reported here was done in collaboration with Pierre L’Ecuyer, Bruno Tuffin, Marvin Nakayama, Michael Gnewuch, Peter Kritzer and Zexin Pan. I received helpful comments on this document from Peter Kritzer, Marvin Nakayama, Pierre L’Ecuyer and two anonymous reviewers.